### v1

@@ 2 DB01068 Clonazepam CTD/ClinicalTrials
3 DB01235 Levodopa DB/CTD/PubChem/ DrugCentral/ClinicalTrials/
4 DB00810 Biperiden DB/CTD/PubChem/ DrugCentral
@@ 6 DB00575 Clonidine ClinicalTrials
7 DB00915 Amantadine DB/PubChem/DrugCentral/ ClinicalTrials
8 DB00989 Rivastigmine DB/CTD/PubChem/ DrugCentral/ClinicalTrials
9 DB00376 Trihexyphenidyl DB/CTD/PubChem/ DrugCentral
@@ 10 DB00215 Citalopram CTD
[…]
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provide reassurance about the reliability of the model.
@@ 4 Conclusion
In the future, we will consider incorporating more biological entities, such as miRNA and proteins, for drug repositioning. Additionally, we will explore more accurate methods to calculate drug (or disease) similarity to reduce data noise.
Key Points We propose DFDRNN, a new deep learning method for drug repositioning, which simultaneously extracts two kinds of features: the similarity feature and the association feature to represent drugs and diseases in each network layer. Based on two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module, DFDRNN captures the potential relationships between the features of drugs and diseases. A cross-dual-domain decoder is proposed to implement a cross-computation between two encoded features to redict drug-disease associations. DFDRNN outperformed the state-of-the-art methods on four benchmark datasets.
5 Acknowledgments

### v2

@@ 2 DB01068 Clonazepam CTD/ClinicalTrials
3 DB01235 Levodopa DB/CTD/DrugCentral/ PubChem/ClinicalTrials
4 DB00810 Biperiden DB/CTD/DrugCentral/ PubChem
@@ 6 DB00575 Clonidine ClinicalTrials
7 DB00915 Amantadine DB/DrugCentral/PubCh-em/ClinicalTrials
8 DB00989 Rivastigmine DB/CTD/DrugCentral/ PubChem/ClinicalTrials
9 DB00376 Trihexyphenidyl DB/CTD/DrugCentral/ PubChem
@@ 10 DB00215 Citalopram CTD
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provides reassurance about the reliability of the model.
@@ 4 Conclusion
For future work, we would like to investigate whether specific enhancements, such as integrating temporal data and handling multi-modal data, could further enhance the performance of DFDRNN. By incorporating temporal data, the model can better understand and predict patterns over time. Integration of data from different modalities could provide a more comprehensive understanding of the input. We aim to identify and implement the appropriate techniques and models for these enhancements. Furthermore, we intend to explore the integration of additional biological entities, such as miRNA and proteins, for drug repositioning.
Acknowledgments